The slope of the line tangent to the graph of у = ln (?) at x = е? is ооо Save Answer
Answer #1

Quer Any y = ln (22) x=p2 Slope of Tangent - dy Here are ((82= (22) = 2- at x = e² a deve o 2 (Option-B) one of the line Tang

Please appreciate my work.
If you have any query regarding this question. You are free to ask.